1. John gave 1/5 of his salary to his parents and spent $240 less than the amount he gave to the parents. He had $1740 left. How much was his salary?
2. A fruit seller sold 1/5 of his pears on Friday and 1/3 of the remainder on Saturday. He found that he had sold 9 fewer pears on Friday than on Saturday. How many pears did He sell altogether?

Answer by mananth(16949)

1. John gave 1/5 of his salary to his parents and spent $240 less than the amount he gave to the parents. He had $1740 left. How much was his salary?

let salary be x
x/5 he gave to parents
He had balance (x-x/5) = 4x/5
He spent (x/5 -240)
So he had balance of 4x/5 -(x/5-240)
4x/5 -(x/5-240)=1740
3x/5 +240 =1740
3x/5 = 1740-240
3x/5 = 1500
3x=5*1500
x= 5*1500/3
x=5*300
=1500
salary was $1500
2. A fruit seller sold 1/5 of his pears on Friday and 1/3 of the remainder on Saturday. He found that he had sold 9 fewer pears on Friday than on Saturday. How many pears did He sell altogether?

Let him have x pears
Friday he sold x/5
balance on Friday = x-x/5 = 4x/5

He sold 1/3 of balance on Saturday
=1/3 * 4x/5= 4x/15 on Saturday
Balance on Saturday = 4x/5 -4x/15= 8x/15

4x/15 -x/5 =9
x/15 =9
x=135
In all he had 135 pears
Plug x in the expression on the number he sold on Friday and Saturday to get the total sold